# Impact of QCD Energy Evolution on Observables in Heavy-Ion Collisions

### v1

@@ I Introduction
Ultra-relativistic heavy-ion collisions are a powerful tool to probe nuclear matter under extreme conditions of temperature and density. Most collision energies at the Relativistic Heavy Ion Collider (RHIC) are high enough to facilitate the transition to the quark-gluon plasma (QGP) Muller:2012zq; Harris:2024aov. Much effort has been devoted to exploring the properties of the QGP, which can lead to a deeper understanding of many-body quantum chromodynamics (QCD). One important insight we have gained is that the matter produced in collisions at RHIC and the Large Hadron Collider (LHC) behaves like a nearly perfect fluid Heinz:2013th; Gale:2013da; Jeon:2015dfa; Romatschke:2017ejr; Heinz:2024jwu, meaning that its viscosity over entropy density ratio is among the lowest realized in nature.
@@
Center-of-mass energies accessible at RHIC and LHC span several orders of magnitude, demanding that the hydrodynamic modeling incorporate a controlled collision energy dependence. This energy dependence is typically embedded within the modeling of the system’s initial conditions. In the case of the Monte Carlo Glauber framework, parameters such as the nucleon-nucleon interaction cross section and an overall normalization factor are tuned to reflect variations in particle production with collision energy Miller:2007ri. In the Eskola-Kajantie-Ruuskanen-Tuominen (EKRT) model Kuha:2024kmq the energy dependence results from the behavior of nuclear parton distribution functions that were fitted to data. Similarly, the IP-Glasma approach Schenke:2012wb, which relies on the color glass condensate (CGC) formalism Iancu:2003xm; Garcia-Montero:2025hys – an effective field theory of QCD – captures energy dependence through adjustments in the saturation scale, guided by the constraints on parameters of the IPSat model Kowalski:2003hm obtained from deep inelastic scattering data Rezaeian:2012ji.
@@
While these approaches have proven effective, successfully reproducing a wide range of observables across different collision energies Schenke:2020mbo, the growing complexity of experimental measurements and the increasing precision of data call for a more refined treatment of the $\sqrt{s}$-dependence in modeling the initial conditions for hydrodynamic simulations.
@@
In this work, and in a companion letter Mantysaari:2025tcg, we expand our analysis of the effects of incorporating the JIMWLK equations Jalilian-Marian:1997qno; Iancu:2000hn; Mueller:2001uk; Lappi:2012vw; Cali:2021tsh into the IP-Glasma model. Including the JIMWLK evolution allows us to describe the nuclei at the appropriate $x$-values, representing the longitudinal momentum fraction of the participating partons, and approximately given by $\langle p_{T}\rangle/\sqrt{s}$, where $\langle p_{T}\rangle$ is the average transverse momentum of produced particles.
@@
In Mantysaari:2025tcg we studied how predicted charged particle multiplicities and transverse momenta in different collision systems were affected by including JIMWLK evolution as compared to the original IP-Glasma model, where the energy dependence follows by adjusting the saturation scale $Q_{s}$ to the appropriate $x$-dependent value. Here, besides providing more details on the combined IP-Glasma+JIMWLK model, we extend our analysis to more differential observables, including various anisotropic flow measurements and correlations between mean transverse momentum and elliptic flow. In particular, we quantify how the evolution in the spatial structure of light and heavy nuclei affects the bulk observables in high-energy collisions.
@@
We note that JIMWLK evolution has been included in heavy-ion initial state descriptions in the past to study the rapidity dependence of the initial geometry Schenke:2016ksl; Schenke:2022mjv; McDonald:2023qwc; Mantysaari:2024qmt. As in Mantysaari:2025tcg, we focus on midrapidity results, and leave an extension to simulations in 3+1D, where the JIMWLK evolution will affect both the collision energy and rapidity dependence of the collisions, to future work.
@@ II IP-Glasma initial state with JIMWLK evolution
To model the spatial configuration of colliding nuclei, initial collision, and pre-equilibrium evolution, we employ the IP-Glasma framework Gale:2012rq; Schenke:2012wb. In its standard formulation—widely used to describe a range of RHIC and LHC data Schenke:2020mbo—the IP-Glasma model encodes the impact parameter $\mathbf{b}_{\perp}$} and Bjorken-$x$ dependence via a local saturation scale $Q_{s}^{2}(x,{\mathbf{b}_{\perp}})$. The saturation scale itself is extracted from the IPSat model Kowalski:2003hm, calibrated against HERA deep inelastic scattering data Rezaeian:2012ji; H1:2009pze.
In contrast to this traditional approach, our current work introduces a dynamically evolving energy dependence by numerically solving the JIMWLK evolution equations. This shift brings a key distinction: while the IP-Sat-based saturation scale yields a static geometric profile with respect to $x$, the JIMWLK evolution captures the growth and smoothing of hadronic and nuclear profiles as $x$ decreases. This results in increasingly diffuse and spatially extended color charge distributions at small-$x$, a behavior observed in prior studies Mantysaari:2022sux; Schlichting:2014ipa, and consistent with the $x$-dependence of generalized parton distribution functions that were constrained with experimental data Moutarde:2018kwr. Furthermore, our approach does not rely on any extracted parametrization of $Q_{s}^{2}$, as we instead solve the perturbative evolution equations for the high-energy nuclear structure.
Our initialization of the small-$x$ evolution uses the same nuclear structure inputs as in earlier versions of the IP-Glasma model Gale:2012rq; Schenke:2012wb; Schenke:2020mbo; Mantysaari:2022sux. At the starting scale $x=x_{0}=0.01$, the color field configuration of each nucleon is encoded in Wilson lines $V({\mathbf{x}_{\perp}})$, defined on a discrete transverse grid. These Wilson lines are obtained from the Yang-Mills equations, using sampled color charges as external currents moving with the speed of light. The color charges are obtained from the McLerran-Venugopalan (MV) model McLerran:1993ni, where they are described as random local Gaussian variables with a correlator
@@
The saturation scale $Q_{s}^{2}$ in the IPsat model depends on the transverse spatial profile of the target, represented by the thickness function $T({\mathbf{b}_{\perp}})$. For nucleons, this profile is modeled as a superposition of $N_{q}$ Gaussian-shaped hot spots Mantysaari:2016ykx; Mantysaari:2020axf; Mantysaari:2025ltq, where each hot spot’s position $\mathbf{b}_{\perp,i}$} is sampled from a Gaussian distribution with width $B_{qc}$. The center of mass of the $N_{q}=3$ hot spots is shifted to the origin after sampling. The resulting expression for the nucleon thickness function is
@@
For nuclear targets, nucleon positions are distributed according to the nuclear density profile and the total thickness function is obtained by summing over the contributions from all constituent nucleons. For large nuclei we employ Woods-Saxon profiles (with a minimum distance of 0.9 fm between the nucleons following Broniowski:2010jd; Schenke:2020mbo), and for smaller nuclei we use results from more sophisticated nuclear structure calculations. The configurations of nucleons for 16O and 20Ne are obtained from ab initio Projected Generator Coordinate Method (PGCM) calculations Giacalone:2024ixe; Giacalone:2024luz, and enforcing two protons and two neutrons to sit close to $\alpha$-cluster centers.  [^1] We note that in Ref. Mantysaari:2025tcg we used the Variational Monte-Carlo (VMC) Carlson:1997qn for the oxygen configurations. The deuteron wave function is sampled from the simple Hulthen wave function with the parameters experimentally determined in Ref. Miller:2007ri.
The model further incorporates event-by-event variations in the local saturation scale, implemented through independent $Q_{s}$ fluctuations for each hot spot. This is done because having such fluctuations is preferred by low-$|t|$ incoherent diffractive $\mathrm{J}/\psi$ production at HERA Mantysaari:2016ykx; Mantysaari:2025idf. Specifically, the normalization of each $T_{q}$ is scaled by a stochastic factor $p_{i}$ drawn from a log-normal distribution
@@
The dependence of Wilson lines on the energy scale, or equivalently on Bjorken-$x$, can be determined by numerically solving the JIMWLK evolution equation Cali:2021tsh; Mueller:2001uk. Adopting the formulation presented in Refs. Lappi:2012vw; Mantysaari:2018zdd, the evolution of the Wilson line $V({\mathbf{x}_{\perp}})$ with rapidity $y=\ln(x_{0}/x)$ can be expressed in a form where updates involve left and right multiplication of $V({\mathbf{x}_{\perp}})$ as follows: [^2] We follow the “square root” running coupling prescription and omit the inclusion of the coupling constant in the noise term, differing slightly from the implementation in Ref. Mantysaari:2018zdd.
@@
Following Refs. Schlichting:2014ipa; Mantysaari:2018zdd, we introduced an infrared regulator $m$ to tame the long-range Coulomb tails that would otherwise cause the proton size to grow faster than allowed by the Froissart bound. This regulator is a free parameter in the model. In the limit $m\to 0$ the above kernel reduces to the perturbative gluon emission kernel $x^{i}/\boldsymbol{x}_{\perp}^{2}$.
@@
with parameters $\mu_{0}=0.28\,\mathrm{GeV}$, $\zeta=0.2$, and $N_{\mathrm{f}}=3$ as in Ref. Mantysaari:2022sux. The scale parameter $\Lambda_{\mathrm{QCD}}$, which controls the running of the coupling in coordinate space, is treated as a free parameter in the model.
@@
    IP-Glasma+JIMWLK: This approach follows the setup of Ref. Mantysaari:2022sux, where the Wilson lines are obtained by solving the full JIMWLK evolution starting from the parametrized initial condition at $x=0.01$ to $x=\langle p_{T}\rangle e^{\pm y}/\sqrt{s_{\mathrm{NN}}}$, with $\langle p_{T}\rangle$ being the charged hadron mean transverse momentum. We use $\left\langle p_{T}\right\rangle=0.5$ GeV at RHIC kinematics and $\left\langle p_{T}\right\rangle=0.65$ GeV at the LHC. The relative variation in $\left\langle p_{T}\right\rangle$ from RHIC to LHC is much smaller than the changes in collision energy when determining the $x$ values, such that the precise choice of $\langle p_{T}\rangle$ values does not play a significant role. In this work, we consider midrapidity kinematics, $y=0$. The model parameters are constrained by fitting to coherent and incoherent $\mathrm{J}/\psi$ photoproduction data from H1, ZEUS, ALICE, CMS, and LHCb, as described in more detail below.
@@
    $Q_{s}(x)$: As a baseline model, we freeze the model parameters in Setup 1 to their values at the initial $x=0.01$. The energy dependence is included by taking the proton saturation scale to be $x$-dependent as described by the IPsat model. This setup is equivalent to the usual IP-Glasma model as used e.g. in Schenke:2020mbo, except that we follow the original implementation of IP-Glasma, with $x$ determined as described above. We do not determine $Q_{s}$ iteratively by solving $x=x({\mathbf{x}_{\perp}})=Q_{s}(x,{\mathbf{x}_{\perp}})/\sqrt{s_{\mathrm{NN}}}$, which has been done e.g. in Schenke:2020mbo.
@@
The non-perturbative model parameters control the proton’s color charge distribution ($c,\tilde{m}$) and its event-by-event geometric fluctuations ($B_{qc},B_{q},\sigma$) at the initial $x=x_{0}$. Furthermore, the coordinate space running coupling scale $\Lambda_{\mathrm{QCD}}$ controls the speed of the JIWMLK evolution. These parameters have been determined in Ref. Mantysaari:2022sux by fitting to a comprehensive set of data on exclusive $\gamma+p\rightarrow\mathrm{J}/\psi+p$ production. This includes measurements from HERA H1:2005dtp; H1:2013okq; ZEUS:2002wfj and from ultraperipheral $p$ + Pb collisions at the LHC Bertulani:2005ru; Klein:2019qfb; ALICE:2014eof; ALICE:2018oyo; LHCb:2014acg; LHCb:2018rcm, spanning a wide range in Bjorken-$x$ from approximately $10^{-2}$ down to $10^{-6}$. The minimum transverse distance between the two hot spots inside a nucleon, $d_{q,\mathrm{min}}$, is set to $0.254$ fm following Ref. Mantysaari:2022ffw. With that, all initial state parameters, except the overall normalization factor for the energy density profile, have been constrained by exclusive vector meson production data.
We also note that recently an updated analysis that also takes into account vector meson photoproduction in $\gamma+\mathrm{Pb}$ collisions, and provides uncertainty estimates for the model parameters, has been presented in Ref. Mantysaari:2025ltq. In the future, it would be beneficial to quantify how uncertainties in the non-perturbative initial state parameters propagate to final state observables in heavy ion collisions.
@@
$m$ $\Lambda_{\rm QCD}$
0.4 GeV 0.04 GeV
Open-source implementations of the IP-Glasma model and the JIMWLK evolution are publicly available ipglasma_code; jimwlk_code. In addition, a modified version of IP-Glasma coupled to JIMWLK evolution can be accessed at ipglasma_jimwlk_code.

### v2

@@ I Introduction
Ultra-relativistic heavy-ion collisions are a powerful tool to probe nuclear matter under extreme conditions of temperature and density. Most collision energies at the Relativistic Heavy Ion Collider (RHIC) are high enough to facilitate the transition to the quark-gluon plasma (QGP) Muller et al. (2012); Harris and Müller (2024). Much effort has been devoted to exploring the properties of the QGP, which can lead to a deeper understanding of many-body quantum chromodynamics (QCD). One important insight we have gained is that the matter produced in collisions at RHIC and the Large Hadron Collider (LHC) behaves like a nearly perfect fluid Heinz and Snellings (2013); Gale et al. (2013a); Jeon and Heinz (2015); Romatschke and Romatschke (2019); Heinz and Schenke (2024), meaning that its viscosity over entropy density ratio is among the lowest realized in nature.
@@
Center-of-mass energies accessible at RHIC and LHC span several orders of magnitude, demanding that the hydrodynamic modeling incorporate a controlled collision energy dependence. This energy dependence is typically embedded within the modeling of the system’s initial conditions. In the case of the Monte Carlo Glauber framework, parameters such as the nucleon-nucleon interaction cross section and an overall normalization factor are tuned to reflect variations in particle production with collision energy Miller et al. (2007). In the Eskola-Kajantie-Ruuskanen-Tuominen (EKRT) model Kuha et al. (2025) the energy dependence results from the behavior of nuclear parton distribution functions that were fitted to data. Similarly, the IP-Glasma approach Schenke et al. (2012), which relies on the color glass condensate (CGC) formalism Iancu and Venugopalan (2003); Garcia-Montero and Schlichting (2025) – an effective field theory of QCD – captures energy dependence through adjustments in the saturation scale, guided by the constraints on parameters of the IPSat model Kowalski and Teaney (2003) obtained from deep inelastic scattering data Rezaeian et al. (2013).
@@
While these approaches have proven effective, successfully reproducing a wide range of observables across different collision energies Schenke et al. (2020), the growing complexity of experimental measurements and the increasing precision of data call for a more refined treatment of the $\sqrt{s}$-dependence in modeling the initial conditions for hydrodynamic simulations.
@@
In this work, and in a companion letter Mäntysaari et al. (2025b), we expand our analysis of the effects of incorporating the JIMWLK equations Jalilian-Marian et al. (1997); Iancu et al. (2001); Mueller (2001); Lappi and Mäntysaari (2013a); Cali et al. (2021) into the IP-Glasma model. Including the JIMWLK evolution allows us to describe the nuclei at the appropriate $x$-values, representing the longitudinal momentum fraction of the participating partons, and approximately given by $\langle p_{T}\rangle/\sqrt{s}$, where $\langle p_{T}\rangle$ is the average transverse momentum of produced particles.
@@
In Mäntysaari et al. (2025b) we studied how predicted charged particle multiplicities and transverse momenta in different collision systems were affected by including JIMWLK evolution as compared to the original IP-Glasma model, where the energy dependence follows by adjusting the saturation scale $Q_{s}$ to the appropriate $x$-dependent value. Here, besides providing more details on the combined IP-Glasma+JIMWLK model, we extend our analysis to more differential observables, including various anisotropic flow measurements and correlations between mean transverse momentum and elliptic flow. In particular, we quantify how the evolution in the spatial structure of light and heavy nuclei affects the bulk observables in high-energy collisions.
@@
We note that JIMWLK evolution has been included in heavy-ion initial state descriptions in the past to study the rapidity dependence of the initial geometry Schenke and Schlichting (2016); Schenke et al. (2022); McDonald et al. (2023); Mäntysaari and Singh (2024). As in Mäntysaari et al. (2025b), we focus on midrapidity results, and leave an extension to simulations in 3+1D, where the JIMWLK evolution will affect both the collision energy and rapidity dependence of the collisions, to future work.
@@ II IP-Glasma initial state with JIMWLK evolution
To model the spatial configuration of colliding nuclei, initial collision, and pre-equilibrium evolution, we employ the IP-Glasma framework Gale et al. (2013b); Schenke et al. (2012). In its standard formulation—widely used to describe a range of RHIC and LHC data Schenke et al. (2020)—the IP-Glasma model encodes the impact parameter $\mathbf{b}_{\perp}$} and Bjorken-$x$ dependence via a local saturation scale $Q_{s}^{2}(x,{\mathbf{b}_{\perp}})$. The saturation scale itself is extracted from the IPSat model Kowalski and Teaney (2003), calibrated against HERA deep inelastic scattering data Rezaeian et al. (2013); Aaron and others (2010).
In contrast to this traditional approach, our current work introduces a dynamically evolving energy dependence by numerically solving the JIMWLK evolution equations. This shift brings a key distinction: while the IP-Sat-based saturation scale yields a static geometric profile with respect to $x$, the JIMWLK evolution captures the growth and smoothing of hadronic and nuclear profiles as $x$ decreases. This results in increasingly diffuse and spatially extended color charge distributions at small-$x$, a behavior observed in prior studies Mäntysaari et al. (2022a); Schlichting and Schenke (2014), and consistent with the $x$-dependence of generalized parton distribution functions that were constrained with experimental data Moutarde et al. (2018). Furthermore, our approach does not rely on any extracted parametrization of $Q_{s}^{2}$, as we instead solve the perturbative evolution equations for the high-energy nuclear structure.
Our initialization of the small-$x$ evolution uses the same nuclear structure inputs as in earlier versions of the IP-Glasma model Gale et al. (2013b); Schenke et al. (2012, 2020); Mäntysaari et al. (2022a). At the starting scale $x=x_{0}=0.01$, the color field configuration of each nucleon is encoded in Wilson lines $V({\mathbf{x}_{\perp}})$, defined on a discrete transverse grid. These Wilson lines are obtained from the Yang-Mills equations, using sampled color charges as external currents moving with the speed of light. The color charges are obtained from the McLerran-Venugopalan (MV) model McLerran and Venugopalan (1994), where they are described as random local Gaussian variables with a correlator
@@
The saturation scale $Q_{s}^{2}$ in the IPsat model depends on the transverse spatial profile of the target, represented by the thickness function $T({\mathbf{b}_{\perp}})$. For nucleons, this profile is modeled as a superposition of $N_{q}$ Gaussian-shaped hot spots Mäntysaari and Schenke (2016); Mäntysaari (2020); Mäntysaari et al. (2025a), where each hot spot’s position $\mathbf{b}_{\perp,i}$} is sampled from a Gaussian distribution with width $B_{qc}$. The center of mass of the $N_{q}=3$ hot spots is shifted to the origin after sampling. The resulting expression for the nucleon thickness function is
@@
For nuclear targets, nucleon positions are distributed according to the nuclear density profile and the total thickness function is obtained by summing over the contributions from all constituent nucleons. For large nuclei we employ Woods-Saxon profiles (with a minimum distance of 0.9 fm between the nucleons following Broniowski and Rybczynski (2010); Schenke et al. (2020)), and for smaller nuclei we use results from more sophisticated nuclear structure calculations. The configurations of nucleons for 16O and 20Ne are obtained from ab initio Projected Generator Coordinate Method (PGCM) calculations Giacalone and others (2025a, b), and enforcing two protons and two neutrons to sit close to $\alpha$-cluster centers.  [^1] We note that in Ref. Mäntysaari et al. (2025b) we used the Variational Monte-Carlo (VMC) Carlson and Schiavilla (1998) for the oxygen configurations. The deuteron wave function is sampled from the simple Hulthen wave function with the parameters experimentally determined in Ref. Miller et al. (2007).
The model further incorporates event-by-event variations in the local saturation scale, implemented through independent $Q_{s}$ fluctuations for each hot spot. This is done because having such fluctuations is preferred by low-$|t|$ incoherent diffractive $\mathrm{J}/\psi$ production at HERA Mäntysaari and Schenke (2016); Mäntysaari and Le (2025). Specifically, the normalization of each $T_{q}$ is scaled by a stochastic factor $p_{i}$ drawn from a log-normal distribution
@@
The dependence of Wilson lines on the energy scale, or equivalently on Bjorken-$x$, can be determined by numerically solving the JIMWLK evolution equation Cali et al. (2021); Mueller (2001). Adopting the formulation presented in Refs. Lappi and Mäntysaari (2013a); Mäntysaari and Schenke (2018), the evolution of the Wilson line $V({\mathbf{x}_{\perp}})$ with rapidity $y=\ln(x_{0}/x)$ can be expressed in a form where updates involve left and right multiplication of $V({\mathbf{x}_{\perp}})$ as follows: [^2] We follow the “square root” running coupling prescription and omit the inclusion of the coupling constant in the noise term, differing slightly from the implementation in Ref. Mäntysaari and Schenke (2018).
@@
Following Refs. Schlichting and Schenke (2014); Mäntysaari and Schenke (2018), we introduced an infrared regulator $m$ to tame the long-range Coulomb tails that would otherwise cause the proton size to grow faster than allowed by the Froissart bound. This regulator is a free parameter in the model. In the limit $m\to 0$ the above kernel reduces to the perturbative gluon emission kernel $x^{i}/\boldsymbol{x}_{\perp}^{2}$.
@@
with parameters $\mu_{0}=0.28\,\mathrm{GeV}$, $\zeta=0.2$, and $N_{\mathrm{f}}=3$ as in Ref. Mäntysaari et al. (2022a). The scale parameter $\Lambda_{\mathrm{QCD}}$, which controls the running of the coupling in coordinate space, is treated as a free parameter in the model.
@@
    IP-Glasma+JIMWLK: This approach follows the setup of Ref. Mäntysaari et al. (2022a), where the Wilson lines are obtained by solving the full JIMWLK evolution starting from the parametrized initial condition at $x=0.01$ to $x=\langle p_{T}\rangle e^{\pm y}/\sqrt{s_{\mathrm{NN}}}$, with $\langle p_{T}\rangle$ being the charged hadron mean transverse momentum. We use $\left\langle p_{T}\right\rangle=0.5$ GeV at RHIC kinematics and $\left\langle p_{T}\right\rangle=0.65$ GeV at the LHC. The relative variation in $\left\langle p_{T}\right\rangle$ from RHIC to LHC is much smaller than the changes in collision energy when determining the $x$ values, such that the precise choice of $\langle p_{T}\rangle$ values does not play a significant role. In this work, we consider midrapidity kinematics, $y=0$. The model parameters are constrained by fitting to coherent and incoherent $\mathrm{J}/\psi$ photoproduction data from H1, ZEUS, ALICE, CMS, and LHCb, as described in more detail below.
@@
    $Q_{s}(x)$: As a baseline model, we freeze the model parameters in Setup 1 to their values at the initial $x=0.01$. The energy dependence is included by taking the proton saturation scale to be $x$-dependent as described by the IPsat model. This setup is equivalent to the usual IP-Glasma model as used e.g. in Schenke et al. (2020), except that we follow the original implementation of IP-Glasma, with $x$ determined as described above. We do not determine $Q_{s}$ iteratively by solving $x=x({\mathbf{x}_{\perp}})=Q_{s}(x,{\mathbf{x}_{\perp}})/\sqrt{s_{\mathrm{NN}}}$, which has been done e.g. in Schenke et al. (2020). This is done to be able to compare more directly with the JIMWLK setup, in which a position dependent $x({\mathbf{x}_{\perp}})$ would require saving the evolution for all $x$ and evaluate different $x$ slices depending on the position. This would require significantly larger computational and memory resources.
@@
The non-perturbative model parameters control the proton’s color charge distribution ($c,\tilde{m}$) and its event-by-event geometric fluctuations ($B_{qc},B_{q},\sigma$) at the initial $x=x_{0}$. Furthermore, the coordinate space running coupling scale $\Lambda_{\mathrm{QCD}}$ controls the speed of the JIWMLK evolution. These parameters have been determined in Ref. Mäntysaari et al. (2022a) by fitting to a comprehensive set of data on exclusive $\gamma+p\rightarrow\mathrm{J}/\psi+p$ production. This includes measurements from HERA Aktas and others (2006); Alexa and others (2013); Chekanov and others (2002) and from ultraperipheral $p$ + Pb collisions at the LHC Bertulani et al. (2005); Klein and Mäntysaari (2019); Abelev and others (2014a); Acharya and others (2019a); Aaij and others (2014, 2018), spanning a wide range in Bjorken-$x$ from approximately $10^{-2}$ down to $10^{-6}$. The minimum transverse distance between the two hot spots inside a nucleon, $d_{q,\mathrm{min}}$, is set to $0.254$ fm following Ref. Mäntysaari et al. (2022b). With that, all initial state parameters, except the overall normalization factor for the energy density profile, have been constrained by exclusive vector meson production data. The values of $\Lambda_{\rm QCD}$ shown in Table 1 are relatively small, which is a result of it being defined in coordinate space (as opposed to the usual momentum space scale) and the fact that we adjust $\Lambda_{\rm QCD}$ directly to control the evolution speed. In comparison, other works (see, e.g., Refs. Albacete et al. (2011); Lappi and Mäntysaari (2013b)) introduce an additional constant $C^{2}$ (which appears in Eq. (7) of Ref. Lappi and Mäntysaari (2013b): $\sim(4/r^{2})C^{2}/\Lambda_{\rm QCD}^{2}$). There, $C^{2}$ is of order 10, effectively achieving the same effect as choosing a smaller $\Lambda_{\rm QCD}$. The need for a small $\Lambda_{\rm QCD}$ originates from the leading order evolution being too fast: In Lappi and Mäntysaari (2016) it was demonstrated that next-to-leading order effects slow down evolution, something we have to capture effectively using the value of $\Lambda_{\rm QCD}$.
We also note that recently an updated analysis that also takes into account vector meson photoproduction in $\gamma+\mathrm{Pb}$ collisions, and provides uncertainty estimates for the model parameters, has been presented in Ref. Mäntysaari et al. (2025a). In the future, it would be beneficial to quantify how uncertainties in the non-perturbative initial state parameters propagate to final state observables in heavy ion collisions.
@@
$m$ $\Lambda_{\rm QCD}$ $K_{\rm JIMWLK}$ $K_{\rm Q_{s}(x),RHIC}$ $K_{\rm Q_{s}(x),LHC}$
0.4 GeV 0.04 GeV 0.115 0.106 0.095
Open-source implementations of the IP-Glasma model and the JIMWLK evolution are publicly available 49; 53. In addition, a modified version of IP-Glasma coupled to JIMWLK evolution can be accessed at 52.